OpenStudy (anonymous):

a body moves along a straight line with a constant acceleration for 20s after it starts from rest it covers a distance s1 in the first 10s and s2 in the next 10s relation between s1 and s2

OpenStudy (anonymous):

ans=s2=3s1

OpenStudy (ajprincess):

s1=1/2a*100 =50a v=10a s2=100a+1/2*a*100 =100a+50a =150a s1/s2=50a/150a s1/s2=1/3 s2=3s1

OpenStudy (anonymous):

thanzzzzz